RHIC Startup: After running good physics stores all night, STAR needed an access when their magnet tripped off..
Vacuum group flashed the TSPs in the Blue 7 o'clock warm straight section at 4:30 yesterday afternoon. They report that there was no observable improvement in the vacuum. PHENIX reports that there was no noticeable improvement in the background rate. So there is no plan to implement a flashing routine.
The store that came up at 17:00 yesterday had high backgrounds from Yellow at PHENIX. MCR had to pull out the collimators (primarily the Yellow vertical collimators) a bit to optimize running.
Stochastic cooling developed a problem It appears it is operating at the wrong gain, making for slow cooling. Mike Brennan found a bad fiber optic connection and "fixed" the problem at about 10:50AM yesterday.
At the afternoon fill, there will be a brief access for repair of the GPIB interface of the Carbon Polarimeter at 12 o'clock, and a repaired card at 2:00 o'clock.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |